# Edge Agentic AI Framework for Autonomous Network Optimisation in O-RAN

### v1

@@ I Introduction
The Open Radio Access Network (O-RAN) architecture enables intelligent network control through RAN Intelligent Controllers (RICs) that support components such as xApps and rApps, which are modular applications that can dynamically manage radio resources. The ability to distribute intelligence in RIC components across RU-edge-cloud continuum allows mobile network operators (MNOs) to optimise performance, leverage cloud elasticity, and meet the diverse demands of different applications and services within the O-RAN framework [2]. Individual components can be invoked manually by the MNO or, more implicitly, through Agentic AI frameworks to meet performance goals [3, 4]. However, integrating LLM-based Agentic AI agents in O-RAN environments presents unique challenges due to task specific operational timescales and control loop requirements. These include ensuring trustworthy decision-making under strict latency and safety constraints, dealing with limited compute and memory resources at the edge, and managing the complexity of large-scale agent coordination.
@@ III System model
Wireless network optimisation faces the ongoing challenge of maintaining optimal Quality of Service (QoS) under constantly changing conditions, including fluctuating traffic loads, interference patterns, and channel fading. Traditional approaches using predefined algorithms struggle to adapt to predictable yet complex scenarios that significantly impact network performance, such as national festivals where large crowds concentrate in limited areas, causing network overload. Our framework addresses these limitations by employing Agentic AI framework for autonomous decision-making that monitors and controls networks in real-time. The proposed framework, illustrated in Fig. 2, leverages a suite of tools to maintain real-time awareness of network conditions and proactively collaborate to predict and mitigate potential outages before they impact user experience. Our system comprises two integrated components operating in a closed-loop fashion:
@@
III-A Realistic Network Modeling
@@
Traffic Dynamics: Gaussian-distributed peaks during morning (7-9 AM, centred at 8:00) and evening (5-8 PM, centred at 18:30) with base load variations, augmented by special event modeling for crowd gatherings during festivals, concerts, and sporting events (up to 5× normal capacity demand).
@@ III-C AI Agent Decision Framework
where $\gamma_{1}$ denotes the critical network threshold, set to 38 dB in this work. The remaining thresholds are defined as follows: $\gamma_{2}$ = 40 dB, $\gamma_{3}$ = 45 dB, and $\gamma_{4}$ = 48 dB.
@@
Step 5 - Performance Feedback: Calculate reward and update system metrics and the model memory.
@@ III-D LSTM xApp Integration into the Agentic AI Framework
Stage 3 – Agent-Driven Interpretation: LLM-based agents interpret LSTM outputs in relation to current network conditions, optimising resource allocation strategies based on capacity, energy constraints, and service-level priorities.
@@ III-E Key Performance Indicators
Outage Rate: Percentage of time SINR falls below 38 dB threshold:
@@
Furthermore, the model features novel persona-based AI agents with specialised roles in which strategic coordinators ensure long-term stability through scalable detection, tactical coordinators enable rapid action response and balance the trade-off between performance and power consumption. Each persona agent employs tailored decision-making strategies and reward mechanisms, allowing the framework to autonomously adapt its anomaly detection and outage prevention based on operational context and service provider priorities.
@@ V Conclusion
Acknowledgment
This research was funded by EPSRC CHEDDAR (EP/X040518/1), UKRI Grant EP/X039161/1.

### v2

@@ I Introduction
The Open Radio Access Network (O-RAN) architecture enables intelligent network control through RAN Intelligent Controllers (RICs) that support components such as xApps and rApps, which are modular applications that can dynamically manage radio resources. O-RAN employs a disaggregated architecture that divides traditional gNodeB functionality across three open units: the Radio Unit (RU), Distributed Unit (DU), and Centralised Unit (CU). This allows mobile network operators (MNOs) to optimise performance, leverage cloud elasticity, and meet the diverse demands of different applications and services within the O-RAN framework [2]. Individual components can be invoked manually by the MNO or, more implicitly, through Agentic AI frameworks to meet performance goals [3, 4]. However, integrating LLM-based Agentic AI agents in O-RAN environments presents unique challenges due to task specific operational timescales and control loop requirements. These include ensuring trustworthy decision-making under strict latency and safety constraints, dealing with limited compute and memory resources at the edge, and managing the complexity of large-scale agent coordination.
@@ III System model
Wireless network optimisation faces the ongoing challenge of maintaining optimal Quality of Service (QoS) under constantly changing conditions, including fluctuating traffic loads, interference patterns, and channel fading. Traditional approaches using predefined algorithms struggle to adapt to predictable yet complex scenarios that significantly impact network performance, such as national festivals where large crowds concentrate in limited areas, causing network overload. Our framework addresses these limitations by employing an Agentic AI framework for autonomous decision-making that monitors and controls networks in real-time. Our novel framework, illustrated in Fig. 2, leverages a suite of tools to maintain real-time awareness of network conditions and proactively collaborate to predict and mitigate potential outages before they impact user experience. Our system comprises two integrated components operating in a closed-loop fashion:
@@
III-A Realistic Network Modelling
@@
Traffic Dynamics: Gaussian-distributed peaks during morning (7-9 AM, centred at 8:00) and evening (5-8 PM, centred at 18:30) with base load variations, augmented by special event modelling for crowd gatherings during festivals, concerts, and sporting events (up to 5× normal capacity demand).
@@ III-B AI Agent Tool Integration
The integration of external APIs in our framework raises potential reliability and security concerns. To mitigate these risks, we implement multi-source data validation by gathering information from different reliable sources and applying a correlation analysis agent to cross-verify data consistency across sources and identify discrepancies before passing it to our LSTM traffic prediction agent.
@@ III-C AI Agent Decision Framework
where $\gamma_{1}$ denotes the critical network threshold, set to 15 dB in this work to meet Ultra-Reliable Low Latency (URLLC) network needs. The remaining thresholds are defined as follows: $\gamma_{2}$ = 18 dB, $\gamma_{3}$ = 20 dB, and $\gamma_{4}$ = 25 dB.
@@
Step 5 - Performance Feedback: Calculate the reward and update the system metrics and model memory.
@@ III-D LSTM xApp Integration into the Agentic AI Framework
Stage 3 – Agent-Driven Interpretation: LLM-based agents interpret the LSTM outputs in relation to current network conditions, optimising resource allocation strategies based on capacity, energy constraints, and service-level priorities.
@@ III-E Key Performance Indicators
Outage Rate: Percentage of time SINR falls below the 15 dB threshold for Ultra-Reliable Low Latency (URLLC):
@@
Furthermore, the model features novel persona-based AI agents with specialised roles in which strategic coordinators ensure long-term stability through scalable detection, tactical coordinators enable rapid action response, and balance the trade-off between performance and power consumption. Each persona agent employs tailored decision-making strategies and reward mechanisms, allowing the framework to autonomously adapt its anomaly detection and outage prevention based on operational context and service provider priorities.